OpenStudy (anonymous):

using L’Hôpital’s Rule evaluate the limit lim x→1 2x^2-x-3/x+1 (fan and medal)

OpenStudy (anonymous):

Do you mean as x goes to -1 ?: \[\frac{2x^{2}-x-3}{x+1}\]

OpenStudy (anonymous):

yes sorry about that @BenjaminEE

OpenStudy (anonymous):

Well since we can't evaluate at -1 when derive the denominator and the nominator. (4x - 1 ) / 1 Now we evaluate this at -1 (4*(-1) - 1) / 1 = -5 / 1 = -5 :)

OpenStudy (anonymous):

@BenjaminEE thank you (: can you help me with another question similar to this one please?

OpenStudy (anonymous):

I can try. But please make a new thread =) It's easier to read.

OpenStudy (anonymous):

okay (: